### v3

# Exact solution of the frustrated Potts model with next-nearest-neighbor interactions in one dimension via AI bootstrapping

Abstract
The one-dimensional (1D) $J_{1}$-$J_{2}$ $q$-state Potts model is solved exactly for arbitrary $q$ by analytically block-diagonalizing the original $q^{2}\times q^{2}$ transfer matrix into a simple $2\times 2$ maximally symmetric subspace, based on using OpenAI’s reasoning model o3-mini-high to exactly solve the $q=3$ case. Furthermore, by matching relevant subspaces, we map the Potts model onto a simpler effective 1D $q$-state Potts model, where $J_{2}$ acts as the nearest-neighbor interaction and $J_{1}$ as an effective magnetic field, nontrivially generalizing a 56-year-old theorem previously limited to the simplest case ($q=2$, the Ising model). Our exact results provide insights to phenomena such as atomic or electronic order stacking in layered materials and the emergence of dome-shaped phases in complex phase diagrams. This work is anticipated to fuel both research in 1D frustrated magnets for recently discovered finite-temperature application potentials and the fast moving topic area of AI in science.
DOI: 10.1103/y5vc-3t6q
I Introduction
Finding novel phases and phase transitions is a central challenge in various research fields, including condensed matter physics, materials science, quantum information, and microelectronics [1]. Unusual phases abound in frustrated magnets [2, 3], which are described typically by the Ising model or the quantum Heisenberg model with competing nearest-neighbor (NN) interaction $J_{1}$ and next-nearest-neighbor (NNN) interaction $J_{2}$ [1, 2, 3, 4].
Another fundamental model of statistical mechanics is the $q$-state Potts model [5, 6, 7, 8], which is a generalization of the Ising model ($q=2$) and can serve as a useful intermediary to study the transition from discrete (Ising) to continuous (Heisenberg) symmetry. In particular, the one-dimensional (1D) $J_{1}$-$J_{2}$ Potts model could be relevant to problems ranging from the out-of-plane stacking of atomic or electronic orders in layered materials, such as charge stripe ordering in La1.67Sr0.33NiO4 [9], the Star-of-David charge-density wave in 1TT-TaS2 [10], and spin spiral ordering in the Weyl semimetal EuAuSb [11]—where in-plane ordering is governed primarily by detailed quantum-mechanical effects, owing to in-plane interactions that are far stronger than those between planes—to a time series with multiple choices at every time step such as table tennis training drill designs.
While the $J_{1}$-$J_{2}$ Ising and Heisenberg models in one dimension [12, 13, 14, 15] and two dimensions [16, 17, 18] have been extensively studied, only the 1D $J_{1}$-$J_{2}$ Ising model has been solved exactly by using the transfer matrix method [19]. Exact solutions of the 1D $J_{1}$-$J_{2}$ Potts model remain unknown as well. The challenge arises from rapid increase in the order of the transfer matrix, which equals $q^{2}$. No wonder a $9\times 9$ matrix for $q=3$ is already hard to solve analytically and diagonalization of a $\left(10^{10^{10}}\right)^{2}\times\left(10^{10^{10}}\right)^{2}$ matrix for $q=10^{10^{10}}$ is simply beyond reach even numerically. Previous studies remarkably reduced the task to numerical calculations for an effective $q\times q$ matrix in the integer-$q$ formalism of the transfer matrix—and for an effective $2\times 2$ matrix in the continuous-$q$ formalism of the transfer matrix where physics is less transparent—however, short of analytic exact results [20]. Hence, an intuitive understanding of the rich phase behaviors in the 1D $J_{1}$-$J_{2}$ Potts model is still lacking. Since the model with $q=3$ already exhibits a distinct ground-state phase behavior from that with $q=2$ (cf. Fig. 2), it is of fundamental importance to exactly solve the model for arbitrary $q$.
Two recent developments shed light on this long-standing problem. The first one is the analytic reduction of the $4\times 4$ transfer matrix for a decorated Ising ladder to an effective $2\times 2$ matrix using symmetry-based block diagonalization, leading to the discovery of spontaneous finite-temperature ultranarrow phase crossover (UNPC), which exponentially approaches the forbidden finite-temperature phase transition in 1D Ising models [21], and the subsequent discovery of in-field UNPC driven by exotic ice-fire states [22, 23, 24]. These findings point out the promising potentials of 1D frustrated magnets in finite-temperature applications; finding exact solutions for 1D frustrated Potts model could define a milestone in this important new direction. The second development is the derivation of an elegant equation, which determines the critical temperature of UNPC in decorated Ising models in an external magnetic field, by OpenAI’s reasoning model o3-mini-high at the first-ever 1000-Scientist AI Jam Session [24]. Hence, the author was inspired to prompt this AI reasoning model progressively and reflectively to handle the transfer matrix in the integer-$q$ formalism for the $q=3$ case—despite quite a few errors in AI’s responses—and eventually have found a symmetry-based block diagonalization that can analytically reduce the $9\times 9$ transfer matrix of the 1D $J_{1}$-$J_{2}$ three-state Potts model to an effective $2\times 2$ matrix (see Supplemental Material I [25]).
For general $q$, the key symmetry is the full permutation symmetry of the $q$ Potts states. In other words, the Hamiltonian (and therefore the transfer matrix in the integer-$q$ formalism) is invariant under any permutation of the labels $\{1,2,3,...,q\}$; its symmetry group is $\mathcal{S}_{q}$. Although the AI failed to go further but warned that the number of permutations increases dramatically as $q$ increases, the analytic results for the $q=2$ and $3$ cases—especially those that both arrive at an effective $2\times 2$ matrix—stimulated the author to realize that since only the largest eigenvalue ($\lambda$) of the transfer matrix matters in the thermodynamic limit, the task is reduced to identify the symmetry-separated subspace that contains $\lambda$. Then, the author found that this subspace is spanned by two maximally symmetric vectors, according to the Perron–Frobenius theorem for positive matrices [26, 27] (see Supplemental Material II). The resulting $2\times 2$ matrix is easy to solve; therefore, the surprisingly simple exact solution of the 1D $J_{1}$-$J_{2}$ Potts model has been obtained for arbitrary $q$.
The maximally symmetric subspace (MSS) method is in spirit similar to the effective Hamiltonian method for studying strongly correlated electronic materials [28, 29, 30]—with the effective model parameters depending on $J_{1}$, $J_{2}$, $q$, and temperature (see Supplemental Material III.A [25]). Furthermore, a rare exact mapping between two canonical Hamiltonians naturally emerges within the MSS method, namely the 1D $J_{1}$-$J_{2}$ $q$-state Potts model can be mapped onto a simpler 1D $q$-state Potts model, where $J_{2}$ acts as the NN interaction and $J_{1}$ as an effective magnetic field. This not only nontrivially generalizes a 56-year-old theorem previously limited to $q=2$ [12] but also profoundly bridges two research areas focused on spontaneous behaviors and external field effects, respectively, thereby demonstrating relevant-subspace-matching as a broadly applicable theoretical approach to tackling complex problems.
Figure 1: Schematics of (a) the single-chain $J_{1}-J_{2}$ Potts model [6] and (b) its equivalent zigzag-ladder representation. The balls stand for the spins with $q$ states. The orange bonds stand for the nearest-neighbor interactions $J_{1}$ and the greenish bonds stand for the next-nearest-neighbor interactions $J_{2}$. The contribution of each bond to the partition function is $x=e^{\beta J_{2}}$, $y=e^{\beta J_{1}/2}$, or $z=e^{\beta J_{1}}=y^{2}$ when its two end spins are the same, and one otherwise.
II Model and Solution
We consider the following Hamiltonian [Fig. 1(a)]
$H=-J_{1}\sum_{i=1}^{2N}\delta({\sigma_{i},\sigma_{i+1}})-J_{2}\sum_{i=1}^{2N}\delta({\sigma_{i},\sigma_{i+2}}),$ (1)
where $\sigma_{i}\in\{1,2,3,...,q\}$ is the spin variable at site $i$. $\delta({\sigma_{i},\sigma_{i+1}})$ is the Kronecker delta (which equals 1 if $\sigma_{i}=\sigma_{i+1}$ and 0 otherwise). $2N$ is the total number of the spins with $\sigma_{2N+1}\equiv\sigma_{1}$ and $\sigma_{2N+2}\equiv\sigma_{2}$, viz., the periodic boundary condition.
To construct the transfer matrix, we use both the overlapping-pairs formulation for Equation (1) with one spin per unit cell to get $\mathbb{T}$ and the equivalent zigzag-ladder version of the model with two spins per unit cell [Fig. 1(b)] to obtain $\mathbb{T}^{\prime}$; they satisfy $\mathbb{T}^{\prime}=\mathbb{T}^{2}$ (see Supplemental Material I [25]). In the thermodynamic limit $N\to\infty$, the partition function $Z=\mathrm{Tr}\,e^{-\beta H}=\lambda^{2N}$, where $\lambda$ is the largest eigenvalue of the transfer matrix $\mathbb{T}$. The free energy per spin is given by
$f=\lim_{N\to\infty}-\frac{1}{2N\beta}\ln Z=-\frac{1}{\beta}\ln\lambda,$ (2)
where $\beta=1/(k_{\mathrm{B}}T)$ with $T$ being the absolute temperature and $k_{\mathrm{B}}$ the Boltzmann constant. $f$ determines physical properties such as the entropy per spin $S=-\partial f/\partial T$, the specific heat $C_{v}=T\partial S/\partial T$, the NN correlation $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=-\partial f/\partial J_{1}$, the NNN correlation $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=-\partial f/\partial J_{2}$, and the energy per spin $E=-\partial f/\partial\beta=-J_{1}\langle\delta(\sigma_{i},\sigma_{i+1})\rangle-J_{2}\langle\delta(\sigma_{i},\sigma_{i+2})\rangle$.
The $\lambda$-containing subspace is spanned by the following two maximally symmetric $q^{2}\times 1$ vectors:
$|\psi_{1}\rangle$ $=$ $\frac{1}{\sqrt{q}}\sum_{s=1}^{q}\sum_{s^{\prime}=1}^{q}\delta(s,s^{\prime})|s,s^{\prime}\rangle,$
$|\psi_{2}\rangle$ $=$ $\frac{1}{\sqrt{q^{2}-q}}\sum_{s=1}^{q}\sum_{s^{\prime}=1}^{q}\left[1-\delta(s,s^{\prime})\right]|s,s^{\prime}\rangle.$ (3)
The resulting transformation matrix $U_{q2}=\{|\psi_{1}\rangle,|\psi_{2}\rangle\}$ is a $q^{2}\times 2$ matrix that projects the $q^{2}\times q^{2}$ transfer matrix $\mathbb{T}$ to the $2\times 2$ block $\mathbb{T}_{2}=U_{q2}^{\top}\,\mathbb{T}\,U_{q2}$, which is decoupled from the rest thanks to different symmetry and given by
$\mathbb{T}_{2}=\left(\begin{array}[]{cc}u&w\\ w&v\\ \end{array}\right)=\left(\begin{array}[]{cc}xy^{2}&\sqrt{q-1}\,y\\ \sqrt{q-1}\,y&x+q-2\\ \end{array}\right)$ (4)
with the shorthand notations $x=e^{\beta J_{2}}$ and $y=e^{\beta J_{1}/2}$. Note that the maximally symmetric subspace means that the expressions of $u$, $v$, and $w$ can be obtained straightforward by combinatorial analysis. The largest eigenvalue of $\mathbb{T}$ is the larger eigenvalue of $\mathbb{T}_{2}$ and is given by
$\lambda=\frac{u+v}{2}+\sqrt{\left(\frac{u-v}{2}\right)^{2}+w^{2}}.$ (5)
III Mapping
The 1D $J_{1}$-$J_{2}$ $q$-state Potts model can be mapped onto a simpler $q$-state Potts model, where $J_{2}$ acts as a NN interaction and $J_{1}$ appears as an effective magnetic field:
$H_{\mathrm{eff}}=-J_{2}\sum_{i=1}^{2N}\delta({\sigma_{i},\sigma_{i+1}})-J_{1}\sum_{i=1}^{2N}\delta({\sigma_{i},1}).$ (6)
Proof: By the Perron–Frobenius theorem [26], the largest eigenvalue of the $q\times q$ transfer matrix of Equation (S5) lives in the MSS spanned by the following two $q\times 1$ vectors:
$|\phi_{1}\rangle$ $=$ $\sum_{s=1}^{q}\delta(s,1)|s\rangle,$
$|\phi_{2}\rangle$ $=$ $\frac{1}{\sqrt{q-1}}\sum_{s=1}^{q}\left[1-\delta(s,1)\right]|s\rangle.$ (7)
The resulting $2\times 2$ matrix [31] is exactly the same as Equation (S6). ∎
This mapping was previously demonstrated for the simplest case by exploiting a special property unique to $q=2$ [12, 13, 14]. Here, its nontrivial generalization to arbitrary $q$ was achieved by matching the MSS of the two $q$-state Potts models. It implies that an antiferromagnetic Potts model in an external magnetic field—which is more tunable in practice—is as fundamental as a Potts model spontaneously frustrated by competing interactions (see Supplemental Material III.B [25]).
With hindsight, if the mapping for arbitrary $q$ had been conjectured, it could have been verified by numerical calculations. Thus, this conjecture would have provided a hint to the general analytical solution.
IV Results and Discussion
The simplicity of Equation (S6) provides an intuitive understanding of rich phase behaviors in the 1D $J_{1}$-$J_{2}$ Potts model. We are interested in frustrated cases where $J_{2}<0$ is antiferromagnetic and have set $J_{2}=1$ as the energy unit from now on. The $q$-dependent phase diagrams given by the normalized entropy $2S(J_{1},T)/\ln q$ are shown in Fig. 2. A $T_{c}$-dome-like phase emerges at low temperature for small $q$ and disappears for large $q$. The phase diagrams given by $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle$ and $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle$ are shown in Fig. 4 in Appendix B.
To get insights to these rich phase diagrams, we start with analyzing ground-state phase behaviors. At $T=0$, the 1D $J_{1}$-$J_{2}$ Potts model for all $q$ values have three phases separated by two critical points (CPs), determined by the relative magnitude of $u$, $w$, $v$ in Equation (S6). The right-hand-side phase for $J_{1}>2$ (where $u$ dominates) is always “ferromagnetic” with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=1$ and $E=-J_{1}-J_{2}$; more precisely, it is a mixture of $q$ ferromagnetic states, i.e., $\frac{1}{\sqrt{q}}(|111...\rangle+|222...\rangle+\cdots+|qqq$$\rangle.$ The “ferromagnetic transition” CP at $J_{1}=2$ has a residual entropy per spin
$S(J_{1}=2,T=0)=\ln\left(\frac{1+\sqrt{4q-3}}{2}\right).$ (8)
The $q=2$ case (i.e., the Ising model) differs from the $q\geq 3$ cases in two aspects: (i) the two CPs for $q=2$ are symmetry related and located at $J_{1}=\pm 2$, while they are located at $J_{1}=0$ and $2$ for $q\geq 3$. (ii) None of the three phases for $q=2$ has a macroscopic degeneracy, while one or two nontrivial states with residual entropy exist for $q\geq 3$. These phases and CPs will be elaborated in passing. Concerning the spin configurations in one dimension, we adopt the convention that the spins are arranged from left to right from now on.
Figure 2: Phase diagrams for $q=2,3,4,\mathrm{and}~10^{6}$. Left panels: Density plots of the normalized entropy $2S(J_{1},T)/\ln q$ in the $J_{1}-T$ plane. Right panels: The temperature dependence of $2S(J_{1},T)/\ln q$ for selected $J_{1}$ values around the critical points. $J_{2}=1$ is set as the energy unit.
$q=2$: The $J_{1}>0$ and $J_{1}<0$ regions are symmetry related by flipping the spins on one sublattice. Consequently, the left-hand-side phase for $J_{1}<-2$ is “antiferromagnetic” with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=0$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=1$, $E=-J_{2}$. The two CPs are located at $J_{1}=\pm 2$. The CP at $J_{1}=-2$ has the same residual entropy as that at $J_{1}=2$ given by Equation (8), i.e., $\ln(\frac{1+\sqrt{5}}{2})$, which is indeed the same as the magnetic-field-induced CP in the 1D Ising model with only NN interactions [32]. The middle phase for $2<J_{1}<2$ is a dimerized state, i.e., a pair of NN spins have the same variable but differ from the last pair, forming the patterns of …11221122… and its three degenerate shifts, leading to $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=1/2$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=0$, and $E=-J_{1}/2$. None of the three phases has a macroscopic degeneracy.
$q\geq 3$: The middle phase for $0<J_{1}<2$ (where $w$ dominates) is a randomly dimerized state (RDS) with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=1/2$, $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=0$, and $E=-J_{1}/2$, where like in the dimerized phase for $q=2$, a pair of NN spins have the same variable but differ from the last pair, so it has q1q-1 choices, yielding a residual entropy of $\ln(q-1)$ per two spins or $\ln\sqrt{q-1}$ per spin. This phase could be relevant to the stacking problem of the Star-of-David charge-density wave in 1TT-TaS2 [10]. The left-hand-side phase for $J_{1}<0$ (where $v$ dominates) is a paramagnet with $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=\langle\delta(\sigma_{i},\sigma_{i+2})\rangle=0$ and $E=0$, where any spin differs from the two last spins, so it has q2q-2 choices, yielding a residual entropy of $\ln(q-2)$ per spin; notably, $q=3$ is special since the residual entropy is zero, forming the patterns of …123123… and its five degenerate states by permutations of {1, 2, 3}, reminiscent of the ABC stacking pattern of the charge stripe order in La1.67Sr0.33NiO4 [9] and the 120 spin spiral order in the Weyl semimetal EuAuSb [11]. In the CP at $J_{1}=0$, any spin differs from its last next-nearest neighbor, so it has q1q-1 choices, yielding a residual entropy of $\ln(q-1)$ per spin; again, $q=3$ is special because its residual entropies at both CPs are equal to $\ln 2$.
Figure 3: The $q\geq 3$ dependence of the zero-temperature normalized entropy $2S(J_{1},0)/\ln q$ for four different regions of $J_{1}$: $S=\ln(q-2)$ for $J_{1}<0$, $S=\ln(q-1)$ for $J_{1}=0$, $S=\ln\sqrt{q-1}$ for $0<J_{1}<2$, and $S=\ln[(1+\sqrt{4q-3})/2]$ for $J_{1}=2$. The $q=2$ case, which need be specially treated for different $J_{1}$ regions (see text), was added for comparison. $J_{2}=-1$.
The $q$ dependence of the residual entropies for the left and middle phases as well as the two CPs is summarized in Fig. 3. For small $q$, the residual entropies of the CPs (dashed lines) are considerably larger than those of their adjacent phases (solid lines). Consequently, each CP develop a V-shape region in the $J_{1}-T$ phase diagram as $T$ increases (Fig. 2, left panels for $q=2,3,4$). The V-shape regions of the two CPs join to create a $T_{c}$-dome-like region for the middle randomly dimerized phase for $q\geq 3$. When placed near the CPs, the system does not follow the common lore—i.e., transition to the adjacent phase with higher macroscopic degeneracy—but transitions to the CP-developed V-shape region, also shown as the flat region in the $T$ curve of the entropy (Fig. 2, right panels for $q=2,3,4$), where the entropy value equals the corresponding CP’s residual entropy.
On the other hand, Fig. 3 reveals that for large $q$, the residual entropies of the CPs (dashed lines) approach those of one of their adjacent phases (solid lines), becoming indistinguishable—no more V-shaped CP regions (Fig. 2, left panel for $q=10^{6}$). When placed near a phase boundary, the system appears to follow the common lore, i.e., transition to the adjacent phase with higher macroscopic degeneracy. In particular, the low-temperature ferromagnetic phase for $J_{1}>2$ would undergo a two-step phase crossover, first to the middle randomly dimerized phase at $k_{\mathrm{B}}T\approx(J_{1}+2J_{2})/\ln q$ and then to the left paramagnetic phase at $k_{\mathrm{B}}T\approx J_{1}/\ln q$.
The $q$-dependent emerging and vanishing of a dome-like structure, controlled by the relative strength of the residual entropies of the phase’s two CPs, provides one more mechanism for forming a dome-shaped phase, a key phenomenon in complex materials, in addition to the two well-known pictures of (i) a preformed order with phase coherence gradually building up [33] and (ii) competing phases [34, 35].
V Summary
The 1D $J_{1}$-$J_{2}$ $q$-state Potts model has been solved exactly by discovering that the largest eigenvalue of the $q^{2}\times q^{2}$ transfer matrix lives in a $2\times 2$ maximally symmetric subspace, which is equivalent to that of the simpler 1D $q$-state Potts model with $J_{2}$ acting as the NN interaction and $J_{1}$ as the magnetic field. The model’s ground state are found to feature three phases separated by two critical points for all $q$ values. The relative strength of the two critical points’ residual entropies is large and small for small and large $q$, respectively, leading to the emerging and vanishing of a dome-shaped randomly dimerized phase for small and large $q$, respectively, providing a new mechanism for forming a dome-shaped phase and insights to the stacking problems of atomic or electronic orders in layered materials. These discoveries were based on using OpenAI’s latest reasoning model o3-mini-high to exactly solve the $q=3$ case, which embodies a new paradigm of data- and value-driven discovery, inspiring scientists to evaluate and connect insights drawn from the vast—though imperfect—information provided by AI.
Note added. We recently employed the MSS method to demonstrate that the ”half ice, half fire” driven ultranarrow phase crossover persists in 1D decorated Potts models for all $q>2$, exhibiting a $T_{0}$-dome structure in the phase diagram that is absent in the $q=2$ case [36].
Figure 4: Phase diagrams in the $J_{1}-T$ plane for $q=2,3,4,\mathrm{and}~10^{6}$ given by the density plots of $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle$ (top panels) and $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle$ (bottom panels). $J_{2}=1$ is set as the energy unit.
Acknowledgment
Brookhaven National Laboratory was supported by U.S. Department of Energy (DOE) Office of Basic Energy Sciences (BES) Division of Materials Sciences and Engineering under contract No. DE-SC0012704.
DATA AVAILABILITY
The data that support the findings of this article are openly available [37].
Appendix A: Software
WOLFRAM MATHEMATICA 14.2 was used to plot Figs. 24 and perform brute-force numerical computation of the eigenvalues of the $q^{2}\times q^{2}$ transfer matrix for $q^{2}$ up to 1024 to verify the exact results. The codes together with that generated by the AI for the $q=3$ case are openly available [37].
VESTA 3.5.8 [38] was used to plot Fig. 1.
Appendix B: Additional phase diagrams
Figure 4 contains additional phase diagrams given by $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle$ and $\langle\delta(\sigma_{i},\sigma_{i+2})\rangle$. The zero-temperature phase transitions as a function of $J_{1}$ are first-order, since $\langle\delta(\sigma_{i},\sigma_{i+1})\rangle=-\partial f/\partial J_{1}$ is discontinuous at both $J_{1}=0$ and $J_{1}=2$; they as a function of negative $J_{2}$ the transitions are continuous at $J_{1}=0$ and first-order at $J_{1}=2$.